OpenStudy (anonymous):

The legs of an isosceles triangle have lengths x + 1 and -x + 7 . The base has length 3x - 3. What is the length of the base? A. 4 B. 6 C. 3 D. Cannot be determined

OpenStudy (anonymous):

ooh lengths nvm

OpenStudy (anonymous):

okay but i am still confused on how you would find this answer

OpenStudy (anonymous):

can you solve \[x+1=-x+7\]?

OpenStudy (anonymous):

in you can't that is fine, i will show you

OpenStudy (anonymous):

no i cant

OpenStudy (anonymous):

a) add \(x\) to both sides, get \[2x+1=7\]

OpenStudy (anonymous):

can you solve that ?

OpenStudy (anonymous):

2x=6 x=3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and the base is \[3x-3\] so put a \(3\) where you see \(x\)

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

3 (3)-3

OpenStudy (anonymous):

yes exactly what is that number?

OpenStudy (anonymous):

9-3 6

OpenStudy (anonymous):

bingo

OpenStudy (anonymous):

so what answer is it

OpenStudy (mathmale):

You have a formula for the length of the base, and you have a value for x. You calculate the length of the base using this info.

OpenStudy (anonymous):

6

OpenStudy (mathmale):

sunshine0829 9-3 6 You obtained that result earlier. Why are you asking "so what answer is it?"